# Approximate independence of permutation mixtures

### v2

@@ 1.4 Organization
The rest of this paper is organized as follows. In Section 2, we review some existing methods of bounding the $\chi^{2}$ divergence between mixture distributions, and show their failure to obtain the tight bound (1) for the toy Gaussian model at the beginning of the introduction. Section 3 and 4 provide the details of the doubly centered expansion and the matrix permanent approach, respectively, where a key inequality in the analysis is a new upper bound on the elementary symmetric polynomial for centered real or complex vectors in Section 3.3. Section 5 provides more discussions and applications of our results, and we defer further discussions on the relationship between the two approaches (Appendix A), proofs of the new inequality (Appendix B) and other main results (Appendix C) to the appendix.
@@ 2 Failure of existing approaches
Before we present our proof of Theorem 1.2, we review several existing approaches to upper bound the statistical distance between mixture distributions. Unfortunately, we will show that all of them fail to yield the bound (1). For these approaches, occasionally we will show the failure under other statistical distances such as the squared TV distance or the KL divergence, both of which are no larger than the $\chi^{2}$ divergence.
@@ 2.1 Reduction to two simple distributions
where the inequality is due to the joint convexity of the KL divergence, $\Pi(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})$ denotes all possible couplings with marginals $\nu_{\mathbb{P}}$ and $\nu_{\mathbb{Q}}$, and $W_{2}$ is the Wasserstein-$2$ distance. For the last term, note that $\theta\mapsto|\sum_{i=1}^{n}\theta_{i}|$ is $1$-Lipschitz under $\|\cdot\|_{2}$, so we can lower bound it as
$W_{2}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})\geq W_{1}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})\geq\mathbb{E}_{\vartheta\sim\nu_{\mathbb{Q}}}\left|\sum_{i=1}^{n}\vartheta_{i}\right|-\mathbb{E}_{\vartheta\sim\nu_{\mathbb{P}}}\left|\sum_{i=1}^{n}\vartheta_{i}\right|=\mu\mathbb{E}\sqrt{\left|2\text{\rm B}\left(n,\frac{1}{2}\right)-n\right|}=\Omega\left(\mu n^{\frac{1}{4}}\right),$
where $\text{\rm B}(\cdot,\cdot)$ denotes the Binomial distribution, and the last step is due to CLT. Consequently, this coupling approach can at best provide an upper bound of $O(\sqrt{n}\mu^{2})$, which not only grows with $n$ but also exhibits a loose dependence on $\mu$ (i.e. $O(\mu^{2})$ instead of $O(\mu^{4})$ for small $\mu$).
@@ 2.2 Reduction to one simple distribution
$\overset{\rm(a)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[W_{2}^{2}\left(\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}\left[{\mathcal{N}}(\vartheta,I_{n})\right],{\mathcal{N}}(\vartheta^{\prime},I_{n})\right)\right]$
$\overset{\rm(b)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[\left\|\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}[\vartheta]-\vartheta^{\prime}\right\|_{2}^{2}\right]\overset{\rm(c)}{\geq}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[\frac{1}{n}\left(\sum_{i=1}^{n}\vartheta_{i}^{\prime}\right)^{2}\right]=\mu^{2},$
@@ 2.3 Method of moments
The previous failures illustrate the importance of directly comparing the mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ rather reducing one or both measures to simple distributions. The method of moments is a powerful approach for performing such a direct comparison. Expanding the likelihood ratio in the orthogonal basis of Hermite polynomials shows that (see, e.g. [HP15, Lemma 2.2] and [WY20, Lemma 9])

### v3

@@ 1.4 Organization
The rest of this paper is organized as follows. In Section˜2, we present several statistical applications of our main theorems, highlighting their usefulness in the analysis of statistical procedures. We then turn to our main results. In Section˜3, we review some existing methods of bounding the $\chi^{2}$ divergence between mixture distributions, and show their failure to obtain the tight bound (1) for the toy Gaussian model at the beginning of the introduction. Section˜4 and 5 provide the details of the doubly centered expansion and the matrix permanent approach, respectively, where a key inequality in the analysis is a new upper bound on the elementary symmetric polynomial for centered real or complex vectors in Section˜4.3. We defer broader discussion of our results (Section˜6), further comments on our proof techniques (Appendix˜A), proofs of the new inequality (Appendix˜B) and other main results (Appendix˜C) to the appendix.
2 Statistical applications
In this section, we present several statistical applications of our main results. These include the identification of the least-favorable prior for the $\ell^{p}$ constrained Gaussian sequence model, an improved differential privacy guarantee for the “shuffled privacy model,” and black-box consistency results for empirical Bayes procedures in compound decision problems. In all three cases, our results yield more general results, with fewer assumptions, than were previously available.
2.1 The least-favorable prior over $\ell^{p}$ balls
Our results give new insights into the classical Gaussian sequence model. Let the observation vector be $X\sim{\mathcal{N}}(\theta,I_{n})$, with an unknown mean vector $\theta\in\mathbb{R}^{n}$ lying in the following $\ell^{p}$ ball $\Theta_{p}(R)$:
Sharp bounds on this minimax risk were first established by Dohono and Johnstone [DJ94], who showed that if $q\geq p\vee 1$, a soft-thresholding estimator with threshold $\mu=\sqrt{2\log(1/R^{p})}$ is asymptotically minimax, with risk $r_{p,q}^{\star}(n,R)=(1+o(1))\mu^{q-p}R^{p}$, provided that the density condition $s:=\frac{nR^{p}}{\mu^{p}}=\omega(1)$ holds. This condition was later removed in [Zha12], who showed by a different argument that the same result holds in the sparse case, when $s=O(1)$.
The most challenging aspect of these results, highlighted in [DJ94], is to obtain a lower bound on the minimax risk $r_{p,q}^{\star}(n,R)$ by bounding the Bayes risk under (a sequence of) carefully chosen priors. Identifying these “least-favorable priors” gives fundamental insight into the statistical structure of the problem, since they capture the core difficulty of the estimation task. Due to the permutation invariance of the parameter space $\Theta_{p}(R)$, it is natural to conjecture that a permutation prior is the asymptotically least favorable, under which the marginal distribution of the observations is a permutation mixture. However, the technical difficulties of analyzing such mixtures meant that [DJ94] used the i.i.d. prior $(1-\alpha)\delta_{0}+\frac{\alpha}{2}(\delta_{-\mu}+\delta_{\mu})$ instead, with $\alpha=(1+o(1))\frac{s}{n}$. [^2] We note for comparison with what follows that their argument applied to the two-point prior $(1-\alpha)\delta_{0}+\alpha\delta_{\mu}$ gives the same lower bound. Of course, this prior is not supported on $\Theta_{p}(R)$; however, the crucial density condition $s=\omega(1)$ is used to guarantee that it approximately concentrates on $\Theta_{p}(R)$ as $n\to\infty$. Nevertheless, their argument does not reveal whether a permutation prior is indeed least favorable.
To analyze the sparse case, Zhang [Zha12] successfully analyzed the permutation prior in the $s=O(1)$ regime. He established that the prior given by $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$, where $v=(\mu,\dots,\mu,0,\dots,0)$ has $s$ nonzero entries, is indeed asymptotically least favorable. The major benefit of this analysis is that it explicitly identifies a bona fide prior on $\Theta_{p}(R)$. However, Zhang’s key step [Zha12, Prop. 1] fails unless $s=O(1)$. What limits the scope of his results is the lack of tools for analyzing the permutation mixture.
To summarize, these results leave open the question of what drives the statistical difficulty of the problem. Is there a fundamental difference between the two regimes $s=\omega(1)$ and $s=O(1)$, or does the same least-favorable prior saturate the minimax risk in both cases? Our results show that the latter is the case.
Lemma 2.1.
Let $q\geq p\vee 1$ and $R=o(1)$. For $\mu=\sqrt{2\log(1/R^{p})}$ and $s=\frac{nR^{p}}{\mu^{p}}=\omega(1)$, the Bayes $\ell^{q}$ risk under the permutation prior $\theta=(v_{\pi(1)},\dots,v_{\pi(n)})$ with $\pi\sim\text{\rm Unif}(S_{n})$ and
is lower bounded by $(1-o(1))(1-c(\varepsilon))\mu^{q-p}R^{p}$, with $c(\varepsilon)\to 0$ as $\varepsilon\to 0^{+}$.
When combined with the analysis in [Zha12] for the $s=O(1)$ regime, Lemma˜2.1 shows upon taking $\varepsilon\to 0^{+}$ that permutation priors are asymptotically least favorable for the minimax risk $r_{p,q}^{\star}(n,R)$ for any choice of $s$.
2.2 Amplification by shuffling for the Gaussian mechanism
Theorem˜1.5 also finds applications to differential privacy [DMNS06], particularly in the context of the shuffled privacy model [EFM+19]. In this model, each client sends her (randomized) message to a secure shuffler, which randomly permutes all incoming messages before forwarding them to the server. This stylized setting captures the behavior of a system in which users’ submissions are stripped of identifying metadata before aggregation, so that an adversary does not know which user is associated with which data point. It is known that random shuffling can strengthen privacy guarantees, a phenomenon that [EFM+19] calls “amplification by shuffling.” Our techniques give a very simple proof of this fact, including for the Gaussian mechanism, for which no such guarantees were previously known.
To this end, consider the standard differential privacy requirement with two neighboring datasets, where the private data of client 1 differs between them. Let $P_{1},\dots,P_{n}$ denote the output distributions of the clients under the first dataset. Under the second dataset, only client 1’s distribution changes, resulting in $(P_{1}^{\prime},P_{2},\dots,P_{n})$. Finally, let $\mathbb{P}_{n}$ and $\mathbb{P}_{n}^{\prime}$ denote the distributions of the shuffled outputs under these datasets. We have the following result.
Lemma 2.2.
Suppose $P_{i}=x_{i}+Z$ and $P_{1}^{\prime}=x_{1}^{\prime}+Z$ for some noise mechanism $Z$, with bounded private data $x_{1},\dots,x_{n},x_{1}^{\prime}\in[0,1]$. For $\varepsilon\in(0,1)$, both the Laplace mechanism $Z\sim\mathrm{Lap}(\frac{1}{\varepsilon})$ and the Gaussian mechanism $Z\sim{\mathcal{N}}(0,\frac{1}{\varepsilon^{2}})$ achieve
${\sf TV}(\mathbb{P}_{n},\mathbb{P}_{n}^{\prime})=O\left(\frac{\varepsilon}{\sqrt{n}}\right).$
Without random shuffling, it is well known that both mechanisms only achieve a TV distance $\Theta(\varepsilon)$ between the output distributions. In other words, random shuffling amplifies privacy. This phenomenon has been studied in [EFM+19, GDD+21, FMT22], where the results require each client’s local mechanism to satisfy $\varepsilon$-LDP (local differential privacy). Our requirement in Lemma˜2.2 recovers this condition for the Laplace mechanism (which is $\varepsilon$-LDP), but our result also applies to the Gaussian mechanism (which does not satisfy $\varepsilon$-LDP for any $\varepsilon>0$), which gives a novel demonstration of amplification by shuffling in this setting.
2.3 Consistency of empirical Bayes procedures for the compound decision problem
Our results can also be applied to obtain general results for the compound decision problem [Rob51]. In this problem, the statistician observes independent data $X_{1},\dots,X_{n}$, where $X_{i}\sim P_{\theta_{i}}$ for $i=1,\dots,n$; the goal is to estimate the vector of parameters under a separable loss $L(\theta,\widehat{\theta}):=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta_{i},\widehat{\theta}_{i})$. This setting is the starting point of the theory of empirical Bayes [Efr19], which views the parameters as having been drawn i.i.d. from a prior distribution, which is in turn estimated from the observations.
An oracle version of the compound decision problem was suggested by Brown and Greenshtein [BG09] (see also [JZ09]), in which the statistician knows the unordered set $\{\theta_{1},\dots,\theta_{n}\}$, but not their correspondence to the observations $(X_{1},\dots,X_{n})$. Equivalently, the observations may be viewed as arising from the permutation mixture $\mathbb{P}_{n}$. As Greenshtein and Ritov [GR09] observe, however, most oracle inequalities in the empirical Bayes literature compare to an oracle that knows $\{\theta_{1},\dots,\theta_{n}\}$ but is restricted to the use of “simple” decision rules, of the form $\widehat{\theta}=(\Delta(X_{1}),\dots,\Delta(X_{n}))$ for some fixed univariate function $\Delta$. Such rules are in fact optimal when the observations arise not from $\mathbb{P}_{n}$ but from its independent counterpart $\mathbb{Q}_{n}$. Denoting by ${\mathcal{D}}^{\mathrm{S}}$ the set of such simple rules, the “Fundamental Theorem of Compound Decisions” (see, e.g., [Cop69]) states that for any $\widehat{\theta}\in{\mathcal{D}}^{\mathrm{S}}$,
$\mathbb{E}[L(\theta,\widehat{\theta})]=\mathbb{E}_{\vartheta\sim\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i}}}\ell(\vartheta,\Delta(X))\,$, (9)
[…]
$\mathbb{E}[L(\theta,\widehat{\theta}^{\mathrm{EB}})]=r^{\mathrm{S}}(\theta)+o(1)\,.$ (10)
For example, Robbins [Rob51] showed that in the Gaussian location model with $\theta_{i}\in\{\pm 1\}$ for $i=1,\dots,n$ and the zero one loss, an empirical Bayes estimator has excess risk $O(n^{-1/2})$ over $r^{\mathrm{S}}(\theta)$. (See [Zha03] for a summary of related results.) In light of (9), such procedures asymptotically match the performance of an oracle that knows the parameters $\theta$ but is restricted to the use of simple rules.
As emphasized by [GR09], however, such guarantees are only partially convincing. Indeed, two important questions remain. In the first place, since $\widehat{\theta}^{\mathrm{EB}}$ is not simple, comparison with the oracle risk $r^{\mathrm{S}}(\theta)$ is not justified. However, since empirical Bayes procedures satisfy the permutation invariance property
$\widehat{\theta}^{\mathrm{EB}}_{\pi(i)}(X_{\pi(1)},\dots,X_{\pi(n)})=\widehat{\theta}^{\mathrm{EB}}_{i}(X_{1},\dots,X_{n}),\quad i\in[n],$
for all $\pi\in S_{n}$, it is more honest to compare against the class ${\mathcal{D}}^{\mathrm{PI}}$ of permutation invariant estimators. This raises the question of characterizing the difference in performance between the best simple decision rule (which is optimal for the i.i.d. setting) and the best permutation-invariant decision rule (which is optimal for the compound decision setting). Moreover, once the restriction to simple estimators is dropped, there is no clear analogue to (9); indeed, in the general setting two natural multivariate Bayesian problems arise (see, e.g. [Wei21]), the first under which the random parameter $(\vartheta_{1},\dots,\vartheta_{n})$ is a uniform random permutation of $(\theta_{1},\dots,\theta_{n})$, and the second under which its coordinates are i.i.d. draws from $\frac{1}{n}\sum_{i=1}^{n}\delta_{\theta_{i}}$. It is the latter problem that appears in (9), but the relationship between the risks of estimators in the two different Bayesian settings is not clear. This raises the question of whether estimators that achieve a good risk bound in one setting also achieve a good risk bound in the other.
Our results allow us to address both concerns. First, Lemma˜2.3 shows that the optimal risks of simple and permutation invariant decision rules agree up to $O(n^{-1/2})$ for general models and bounded losses. Denote the optimal risk among permutation invariant decision rules by ${\mathcal{D}}^{\mathrm{PI}}$:
[…]
Theorem˜1.5 implies the following.
Lemma 2.3.
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq M\sqrt{6{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(e\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))^{3{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})}/n}.$
Lemma˜2.3 shows that for a large class of models and losses, $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(n^{-1/2})$. This strengthens a “folklore” result in [HR55] that shows that the gap is $o(1)$. Combined with (10), these results yield honest oracle inequalities for empirical Bayes procedures. Our improvement is salient for the majority of parametric empirical Bayes problems (like that of Robbins) for which the excess risk over simple decision rules is $O(n^{-1/2})$.
The connection between Lemma˜2.3 and Theorem˜1.5 is the Bayesian perspective alluded to above. Both $r^{\mathrm{S}}(\theta)$ and $r^{\mathrm{PI}}(\theta)$ can be analyzed within the Bayesian framework under which $\vartheta$ is a uniform random permutation of $\theta$. In this formulation, comparing $r^{\mathrm{S}}(\theta)$ with $r^{\mathrm{PI}}(\theta)$ reduces to comparing the conditional distributions $P_{\vartheta_{i}\mid X_{i}}$ and $P_{\vartheta_{i}\mid X^{n}}$, where Theorem˜1.5 plays a central role.
Second, Lemma˜2.4 gives a general comparison between the two different Bayesian settings described above for arbitrary estimators. In particular, it implies that for general bounded losses, estimators that have vanishing risk under the i.i.d. prior also have vanishing risk under the more complicated permutation prior.
Lemma 2.4.
For any estimator $\widehat{\vartheta}=\widehat{\vartheta}(X)$ and separable loss $L(\vartheta,\widehat{\vartheta})=\frac{1}{n}\sum_{i=1}^{n}\ell(\vartheta_{i},\widehat{\vartheta}_{i})$, we have
$|\mathbb{E}_{\mathbb{P}}[L(\vartheta,\widehat{\vartheta})]-\mathbb{E}_{\mathbb{Q}}[L(\vartheta,\widehat{\vartheta})]|\leq\sqrt{e(\chi^{2}({\mathcal{P}})+1)}\cdot\frac{1}{n}\sum_{i=1}^{n}\sqrt{\mathrm{Var}_{\mathbb{Q}}[\ell(\vartheta_{i},\widehat{\vartheta}_{i})]},$ (11)
where $\chi^{2}({\mathcal{P}})$ is any upper bound in Theorem˜1.2 for the distribution family ${\mathcal{P}}=\{P_{\theta_{1}},\dots,P_{\theta_{n}}\}$, and $\mathbb{P}$ and $\mathbb{Q}$ denote the joint distributions of $(\vartheta,X)$ under the permutation prior and i.i.d. prior, respectively.
Moreover, if $0\leq\ell(\cdot,\cdot)\leq M$, then
$\mathbb{E}_{\mathbb{P}}[L(\vartheta,\widehat{\vartheta})]\leq\mathbb{E}_{\mathbb{Q}}[L(\vartheta,\widehat{\vartheta})]+\sqrt{eM(\chi^{2}({\mathcal{P}})+1)\mathbb{E}_{\mathbb{Q}}[L(\vartheta,\widehat{\vartheta})]}\,$. (12)
In particular, if $M(\chi^{2}({\mathcal{P}})+1)$ is bounded, then $\mathbb{E}_{\mathbb{Q}}[L(\vartheta,\widehat{\vartheta})]\to 0$ implies $\mathbb{E}_{\mathbb{P}}[L(\vartheta,\widehat{\vartheta})]\to 0$.
We note that further improvements to Lemma˜2.3 are possible for specific losses, via more specialized arguments. For example, an alternative strategy is developed in [GR09] for the quadratic loss $L(\theta,\widehat{\theta})=\frac{1}{n}\|\theta-\widehat{\theta}\|_{2}^{2}$. For several classes $\mathcal{P}$} (such as the Gaussian location model), [GR09] shows that $r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)=O_{{\mathcal{P}}}(n^{-1})$ under quadratic loss, improving upon the $O_{{\mathcal{P}}}(n^{-1/2})$ upper bound in Lemma˜2.3, which holds for general losses. An adaptation of their proof technique can be used to extend this result to broader classes $\mathcal{P}$} under weaker assumptions.
Lemma 2.5.
Let $P_{\theta_{1}},\dots,P_{\theta_{n}}\in{\mathcal{P}}$ with ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})<\infty$, and $|\theta_{i}|\leq M$ for all $i\in[n]$. Then under the quadratic loss $L(\theta,\widehat{\theta})=\frac{1}{n}\|\theta-\widehat{\theta}\|_{2}^{2}$,
$r^{\mathrm{S}}(\theta)-r^{\mathrm{PI}}(\theta)\leq\frac{4M^{2}{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})(1+{\mathsf{C}}_{\chi^{2}}({\mathcal{P}}))}{n}.$
We remark that the proof of Lemma˜2.5 follows a different strategy than the other arguments in this paper. The key step in the proof of Lemma˜2.5 is to apply convexity with a clever coupling of [GR09] to obtain a Hellinger upper bound between two permutation mixtures, as stated in Lemma˜C.2; however, as witnessed in Section˜3, this convexity-based idea is insufficient to establish (1). Moreover, even for the quadratic loss, Lemma˜2.5 is still not optimal: in our follow-up work [HNWSW25], we use the techniques of the present paper to obtain the first major improvement of Lemma˜2.5 for Gaussian location models since [GR09]; see [HNWSW25, Theorem 4.1].
@@ 3 Failure of existing approaches
Before we present our proof of Theorem˜1.2, we review several existing approaches to upper bound the statistical distance between mixture distributions. Unfortunately, we will show that all of them fail to yield the bound (1). For these approaches, occasionally we will show the failure under other statistical distances such as the squared TV distance or the KL divergence, both of which are no larger than the $\chi^{2}$ divergence.
@@ 3.1 Reduction to two simple distributions
where the inequality is due to the joint convexity of the KL divergence, $\Pi(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})$ denotes all possible couplings with marginals $\nu_{\mathbb{P}}$ and $\nu_{\mathbb{Q}}$, and $W_{2}$ is the Wasserstein-$2$ distance. For the last term, since $\nu_{\mathbb{P}}$ is supported on the set $L:=\{\theta\in\{\pm\mu\}^{n}:\sum_{i=1}^{n}\theta_{i}=0\}$, we can lower bound it as
$W_{2}^{2}(\nu_{\mathbb{P}},\nu_{\mathbb{Q}})\geq\mathbb{E}_{\vartheta\sim\nu_{\mathbb{Q}}}\left[\min_{\theta_{0}\in L}\|\vartheta-\theta_{0}\|_{2}^{2}\right]=\mathbb{E}_{\vartheta\sim\nu_{\mathbb{Q}}}\Bigg{[}4\mu^{2}\Big{|}\sum_{i=1}^{n}\mathbbm{1}_{\vartheta_{i}=\mu}-\frac{n}{2}\Big{|}\Bigg{]}=\Omega(\sqrt{n}\mu^{2}),$
where $\sum_{i=1}^{n}\mathbbm{1}_{\vartheta_{i}=\mu}$ follows a binomial distribution $\mathsf{B}(n,\frac{1}{2})$ under $\nu_{\mathbb{Q}}$, and the last step is due to the CLT. Consequently, this coupling can at best provide an upper bound of $O(\sqrt{n}\mu^{2})$, which not only grows with $n$ but also exhibits a loose dependence on $\mu$ (i.e. $O(\mu^{2})$ instead of $O(\mu^{4})$ for small $\mu$).
@@ 3.2 Reduction to one simple distribution
$\overset{\rm(a)}{\geq}\frac{1}{2}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[W_{2}^{2}\left(\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}\left[{\mathcal{N}}(\vartheta,I_{n})\right],{\mathcal{N}}(\vartheta^{\prime},I_{n})\right)\right]$
$\overset{\rm(b)}{\geq}\frac{1}{2}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\left[\left\|\mathbb{E}_{\vartheta\sim\nu_{\vartheta^{\prime}}}[\vartheta]-\vartheta^{\prime}\right\|_{2}^{2}\right]\overset{\rm(c)}{\geq}\frac{1}{2}\mathbb{E}_{\vartheta^{\prime}\sim\nu_{\mathbb{Q}}}\Bigg{[}\frac{1}{n}\Bigg{(}\sum_{i=1}^{n}\vartheta_{i}^{\prime}\Bigg{)}^{2}\Bigg{]}=\frac{\mu^{2}}{2},$
@@ 3.3 Method of moments
The previous failures illustrate the importance of directly comparing the mixtures $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$ rather than reducing one or both measures to simple distributions. The method of moments is a powerful approach for performing such a direct comparison. Expanding the likelihood ratio in the orthogonal basis of Hermite polynomials shows that (see, e.g. [HP15, Lemma 2.2] and [WY20, Lemma 9])